OpenStudy (sasogeek):

is it okay to do this...? and did i get it correct anyway? the third term of an exponential sequence is 12 and the 9th term is 768. find the values of the common ratio. \[\large ar^2=12 \]\[\large ar^8=768\] \[\large \frac{ar^8}{ar^2}=\frac{768}{12}\] \[\large r^6=64 \] \[\large r=2 \ , \ r=-2\] I forgot the formulas and decided to go with this but i don't know if it is correct or even if it's acceptable... can anyone confirm for me please?

OpenStudy (sasogeek):

truth is i've even forgotten if this is exactly how to do it lol. i literally have no memory of the classes i took for this topic

sam (.sam.):

yeah you can do that

Parth (parthkohli):

You are perfectly correct, math can be done in infinite ways :)

OpenStudy (mimi_x3):

it looks like a geometric progression.. i dont know what's an exponential sequence.. http://en.wikipedia.org/wiki/Exponential_sheaf_sequence

sam (.sam.):

\[a_3=12\] \[a_9=768\] ------------- \[ar^2=12\] \[ar^8=768\] \[\frac{12}{r^2}r^8=768\] \[r^6=64\] \[r= \pm 2\]

OpenStudy (sasogeek):

same thing :) different name

sam (.sam.):

GP formula \[\Huge T_n=a_1r^{n-1}\]

OpenStudy (sasogeek):

thanks :)

sam (.sam.):

np